OpenStudy (anonymous):

Find the arc length of the graph of the function 3*sqrt(x^3) from x=5 to x=8

OpenStudy (anonymous):

use \[\int_a^b \sqrt{1+(f'(x))^2}dx\]

OpenStudy (anonymous):

the derivative is \[f'(x)=\frac{9\sqrt{x}}{2}\] so when you square it you get \[\frac{81x}{4}\]and then you want \[\int_5^8 \sqrt{1+\frac{81x}{4}}dx\][